# Detecting Visual Information Manipulation Attacks in Augmented Reality: A Multimodal Semantic Reasoning Approach

### v1

@@ Abstract
The virtual content in augmented reality (AR) can introduce misleading or harmful information, leading to semantic misunderstandings or user errors. In this work, we focus on visual information manipulation (VIM) attacks in AR where virtual content changes the meaning of real-world scenes in subtle but impactful ways. We introduce a taxonomy that categorizes these attacks into three formats: character, phrase, and pattern manipulation, and three purposes: information replacement, information obfuscation, and extra wrong information. Based on the taxonomy, we construct a dataset, AR-VIM. It consists of 452 raw-AR video pairs spanning 202 different scenes, each simulating a real-world AR scenario. To detect such attacks, we propose a multimodal semantic reasoning framework, VIM-Sense. It combines the language and visual understanding capabilities of vision-language models (VLMs) with optical character recognition (OCR)-based textual analysis. VIM-Sense achieves an attack detection accuracy of 88.94% on AR-VIM, consistently outperforming vision-only and text-only baselines. The system reaches an average attack detection latency of 7.07 seconds in a simulated video processing framework and 7.17 seconds in a real-world evaluation conducted on a mobile Android AR application.
@@ 1 Introduction
Augmented Reality (AR) overlays virtual content onto the real world to enhance user perception, improve decision-making, and enrich interactive experiences. While AR offers benefits across domains such as education, navigation, and entertainment, it also introduces new risks when virtual content is poorly designed or intentionally misleading, creating visual attacks in AR[13, 32, 31, 44, 58]. In particular, visual information manipulation attacks, in which virtual content changes the meaning or interpretation of real-world scenes, can lead users to serious consequences, such as misunderstanding warnings, misidentifying important objects, or taking action based on incorrect, manipulated information[47, 43, 50]. For example, in Fig.˜1, the virtual object (the horizontally flipped “3") is overlaid on a real-world object, subtly altering the exit number from “3" to “8." While visually similar, the change has significant semantic implications and may mislead users during navigation. These attacks are particularly critical because they are often subtle, visually realistic, and seamlessly integrated into the scene, making them difficult to be noticed by users, or detected with automated evaluation tools.
@@
Detecting visual information manipulation attacks requires reasoning about both the virtual content and how it interacts with the real-world scene. Prior work on AR content quality assessment has primarily focused on evaluating visual properties of virtual elements, such as shadow consistency[3], lighting effect[24], and spacial alignment[27, 35, 54, 6] to determine how well the virtual content blends into the environment, or whether it affects users’ perception[52, 38, 18]. While important for visual coherence, these approaches fall short when it comes to semantic-level threats, where content may appear realistic and reasonable, yet alter the meaning of the scene. In such cases, virtual elements may adopt a consistent visual style while subtly introducing misinformation that leads to misunderstanding or misuse. Other work has focused more directly on safety concerns in AR, such as in the context of obstruction attacks[13, 31, 20, 58], where virtual content blocks important real-world information and potentially lead to accidents or task failures. However, while obstruction attacks involve physical concealment, they fail to capture scenarios where virtual content is non-obstructive or only partially overlaps with key information, yet still alters the semantic meaning of the scene. For example, in Fig.˜2, additional virtual guidance is inserted into the scene, directing the user toward an incorrect action without blocking any critical real-world element.
@@
Beyond AR-specific studies, there has been numerous studies in computer vision and multimodal machine learning that focused on scene understanding, including semantic segmentation[40, 15], image captioning[51, 26] and visual question answering (VQA)[5, 53]. These methods demonstrate impressive capabilities in extracting semantic meaning from complex visual inputs. However, they are typically developed and evaluated outside the AR context. As a result, these models’ ability to analyze the presence or potential impact of augmented elements remains underexplored. In AR scenarios, this ability to analyze virtual content is important: while users can only see the augmented view, a robust detection system should evaluate both the raw scene and its augmented counterpart to determine whether virtual content has introduced detrimental changes.
@@
These challenges call for new approaches to evaluate AR content, which go beyond visual realism or obstruction, and instead focus on the semantic impact of virtual elements. In particular, detecting whether virtual content manipulates real-world information requires understanding not just what objects are present, but how their meaning changes when combined. Recent advances in vision-language models (VLMs) with generative capabilities[42, 21, 4] have demonstrated strong capabilities in multimodal reasoning, including fine-grained scene interpretation and textual analysis. Moreover, VLMs have been proven effective in evaluating artificial content[23] and sensitive to virtual content in AR[19]. These capabilities make VLMs well-suited for assessing whether AR content introduces semantic inconsistencies or misleading information.
@@
We introduce AR-VIM, a dataset of 452 raw-AR video pairs collected in 202 different real-world scenes. The videos are annotated to reflect various attack types and semantic distortions. To validate the dataset’s labels, we conducted an Institutional Review Board (IRB)-approved user-based data labeling task. The results demonstrate that our labeling aligns closely with human perception. The dataset is available on GitHub.111https://github.com/YM-Xiu/AR-VIM
We develop a multimodal semantic reasoning framework, VIM-Sense, that combines VLMs with OCR-based text extraction to detect and analyze semantic-level information manipulation attacks in AR scenes. On AR-VIM, our system reaches an attack detection accuracy of 88.94% and a detection latency of 7.07 s. We also evaluate our system on a real-world Android AR application, achieving a detection latency of 7.17 s.
The remaining sections of this paper are organized as follows: Section˜2 reviews prior work, followed by Section˜3 which introduces our taxonomy of visual information manipulation attacks. Section˜4 describes the construction and annotation of the AR-VIM dataset. Section˜5 presents the design and implementation of our system, VIM-Sense. In Section˜6, we evaluate the VIM-Sense on both the AR-VIM dataset and a real-world AR application. Finally we discuss the limitations and directions for future work in Section˜7 and conclude the paper in Section˜8.
@@ 2.1 AR Virtual Content Assessment
Augmented reality research has long examined how virtual content can be designed and evaluated to ensure a pleasant user experience. A significant body of work focuses on assessing the visual properties of virtual elements to determine whether they blend seamlessly into the real world. Kim et al.[27] investigated artifacts of virtual content that arise when physics simulations are used to model interactions between real and virtual objects in AR. Li et al. [35] proposed a method for evaluating misalignment between real and virtual objects in a head-mounted AR assembly assistance system using depth map comparison. Golyadkin et al.[22] proposed a data synthesis and benchmarking framework for manga colorization, which incorporates image inpainting techniques to support consistent colorization of sequential pages and assess the inpainting quality. On the user perception side, Wang et al.[52] conducted a user study to investigate how texture density, luminance contrast, and color contrast of virtual objects affect depth perception in optical see-through AR. Liu et al.[38] introduced the concept of visual confusion as a type of distortion affecting perceptual quality when both real and virtual contents contain complex textures. Duan et al.[18] proposed a framework for evaluating perceptual quality in AR scenes based on visual confusion theory. While these works provide insights into visual realism, perceptual quality, and alignment, they primarily focus on how well virtual content fits into the physical scene or how it affects users’ visual perception. They do not address whether the virtual content introduces semantic distortions or changes the meaning of the scene.
Aside from perceptual quality, some prior works have explored how virtual content can interfere with users’ access to real-world information. Cheng et al.[13] conduct a user study to investigate how people perceive and react to visual obstruction attacks in AR. Lebeck et al. proposed Arya[31, 32], which detects obstructions by using system sensors to identify predefined safety-critical real-world objects. It ensures that critical items are not obstructed by modifying virtual content based on predefined policies. Xiu et al. proposed VIDDAR[58, 57], which integrates generative AI models with object detection models to identify critical real-world objects and determine if a key object is obstructed or made to be visually misleading. These approaches take scene information and user safety into consideration when assessing the virtual content. However, they share a common assumption that safety threats in AR arise when important real-world objects are physically occluded. Consequently, they cannot handle more subtle cases of visual information manipulation, where virtual content does not fully block real-world objects but instead modifies or replaces their meaning.
@@ 2.2 Scene Understanding
Recently, numerous studies in computer vision that focused on scene understanding, which aims to extract semantic meaning from images through tasks such as semantic segmentation, image captioning, and visual question answering (VQA). Long et al.[40] proposed fully convolutional networks for semantic segmentation, which assigns a categorical label to each pixel in an image. Vinyals et al.[51] proposed an image captioning system that uses deep neural networks to extract image features and generate natural language descriptions for the image. Antol et al.[5] proposed the task of VQA, which requires answering open-ended questions about images using natural language. These early works laid the foundation for a wide range of scene understanding tasks. More recently, the field has seen a shift toward foundation models. Kirillov et al.[28] proposed the Segment Anything Model (SAM), a foundation model for semantic image segmentation that supports prompt-based interaction. Generative models with visual modality, also known as VLMs, such as GPT-4V[42], Gemini[21], and Claude[4], have shown superior performance in some scene understanding benchmarks[12, 5, 2, 25]. Overall, recent methods have demonstrated strong capabilities in understanding scene content and identifying potentially harmful or misleading information. However, most of these studies were conducted and evaluated using standard, general-purpose vision-language benchmarks. As a result, their potential for understanding virtual content in AR, where it is essential to analyze both the real-world scene and its augmented counterpart, remains largely underexplored.
@@ 2.3 VLMs for Semantic Reasoning in AR Settings
As mentioned above, the ability of scene understanding methods to operate effectively in AR settings remains unclear. Recently, efforts have begun to explore the use of VLMs for reasoning about artificial content, automatically generated content and virtual content in AR, especially in the context of safety and usability. Guo et al.[23] proposed UGCG-GUARD, a content moderation system for detecting image-based promotions of unsafe user-generated content, leveraging VLMs to perform contextual detection. Srinidhi et al. proposed XaiR[48], a platform that integrates VLMs with XR headsets, which enables real-time semantic reasoning and AR content placement. Sharma et al. proposed OCTO+[45], a framework for automatic open-vocabulary object placement in AR, which leverages VLMs to reason about spatial semantics and place virtual objects in valid regions. Duan et al.[19] evaluated several state-of-the-art VLMs for understanding and describing AR-generated scenes, highlighting that VLMs excel at detecting and assessing obvious virtual content. Nonetheless, these work did not dive into the visual safety concerns in AR settings. In this work, we aim to fill this gap by combining VLM-based semantic reasoning with OCR-based textual grounding to detect visual information manipulation attacks in AR scenes.
@@ 3.1 Attack Format
Character Manipulation: Character manipulation refers to attacks where individual characters in a real-world scene are modified through virtual overlays. These changes may appear subtle (e.g., “3" to “8" or “O" to “Q"), but can significantly alter the meaning of textual information, such as room numbers, prices, or signs.
Denote as the ordered sequence of characters in , and as the ordered sequence of characters in . These two sequences have exactly the same length, which means the characters are only manipulated, while not erased or added. Therefore, the character in is manipulated if:
@@
Pattern Manipulation: Pattern manipulation refers to non-textual visual elements in the scene such as colors, symbols, or directional cues that are altered through virtual content in a way that changes the scene’s semantic meaning. Unlike character or phrase manipulation, which operates on textual information, pattern manipulation affects visual structures that convey meaning independently of text.
@@ 3.2 Attack Purpose
This condition indicates that the original information is no longer present in the AR image, and has been replaced by another token that conveys a different semantic meaning. The substitution may occur in either the textual domain () or the visual domain (), as long as the replacement alters the perceived meaning of the original scene content.
@@
Extra Wrong Information: Extra wrong information attack is a case where new virtual content introduces semantic elements into the AR image that were not present in the raw image and that convey misleading or incorrect meanings. Unlike information replacement or obfuscation, this type of attack does not interfere with existing real-world information but instead adds false signals that may confuse the user or alter the interpretation of the scene. An additional wrong information attack is said to occur if:
@@ 3.3 Attack Taxonomy Summary
Based on the definitions above, we categorize visual information manipulation attacks in AR by combining three attack formats with three attack purposes. This gives a conceptual space of possible attack types. However, not all combinations are applicable in real-world AR scenarios. In particular, we only consider character manipulation in the context of information replacement, since manipulating characters at the level of obfuscation or false addition is rare or ill-defined. For example, it is difficult to obfuscate a character without affecting its surrounding word or phrase, and adding an isolated character into the scene without meaningful context rarely conveys independent semantic intent. The resulting taxonomy consists of seven valid attack types, as shown in Table˜1. Each type corresponds to a unique way that virtual content can semantically alter the scene, whether through subtle text edits or high-level visual modifications. Some examples of the attack are shown in Fig.˜3.
@@ 4.1 VIM Attack Design Methodology
Semantic Alteration with Minimal Virtual Content Insertion: For each scene, our objective was to introduce the minimal amount of virtual content required to alter the scene’s meaning. For example, in character replacement, we selected individual characters whose modification would flip the semantic interpretation (e.g., “3” to “8”), while leaving the surrounding context unchanged. This ensures the manipulation is subtle yet capable of misleading the user.
Visual Style Coherence: To avoid triggering user attention, we ensured that the visual style of the virtual content closely matched the real-world context, which includes color style, lighting effect, and spatial alignment. This is especially important for information replacement and information obfuscation attacks, where the effectiveness of the manipulation relies on the user’s inability to recognize the virtual content. For extra wrong information attacks, we focused on adding semantically misleading but visually reasonable content.
Attack-Non-Attack Pairing Strategy: To ensure controlled comparisons and reliable evaluation of VIM detection, we designed the corresponding non-attack case for most of the attack cases. For information obfuscation, we created a non-attack version by repositioning the virtual content to a different part of the scene that does not obfuscate the key semantic elements in the real world. For information replacement and extra wrong information attacks, we generated visually similar virtual content in terms of color, style, and size, but with either misaligned placement or semantically non-adversarial meaning. For example, in Fig.˜3. (b), the attack shows a misleading phrase like “Free Parking” and the non-attack version might show “Parking Prohibited” in the same font and location, preserving visual realism while avoiding deception.

### v3

@@ Abstract
The virtual content in augmented reality (AR) can introduce misleading or harmful information, leading to semantic misunderstandings or user errors. In this work, we focus on visual information manipulation (VIM) attacks in AR, where virtual content changes the meaning of real-world scenes in subtle but impactful ways. We introduce a taxonomy that categorizes these attacks into three formats: character, phrase, and pattern manipulation, and three purposes: information replacement, information obfuscation, and extra wrong information. Based on the taxonomy, we construct a dataset, AR-VIM, which consists of 452 raw-AR video pairs spanning 202 different scenes, each simulating a real-world AR scenario. To detect the attacks in the dataset, we propose a multimodal semantic reasoning framework, VIM-Sense. It combines the language and visual understanding capabilities of vision-language models (VLMs) with optical character recognition (OCR)-based textual analysis. VIM-Sense achieves an attack detection accuracy of 88.94% on AR-VIM, consistently outperforming vision-only and text-only baselines. The system achieves an average attack detection latency of 7.07 seconds in a simulated video processing framework and 7.17 seconds in a real-world evaluation conducted on a mobile Android AR application.
@@ Introduction
Augmented Reality (AR) overlays virtual content onto the real world to enhance user perception, improve decision-making, and enrich interactive experiences. While AR offers benefits across domains such as education, navigation, and entertainment, it also introduces new risks when virtual content is poorly designed or intentionally misleading, creating visual attacks in AR[13, 33, 32, 45, 59]. In particular, visual information manipulation attacks, in which virtual content changes the meaning or interpretation of real-world scenes, can lead to serious consequences, such as misunderstanding warnings, misidentifying important objects, or taking action based on incorrect, manipulated information[48, 44, 51]. For example, in Fig.˜1, the virtual object (the horizontally flipped "3") is overlaid on a real-world object, subtly altering the exit number from "3" to "8." While visually similar, the change has significant semantic implications and may mislead users during navigation. These attacks are particularly critical because they are often subtle, visually realistic, and seamlessly integrated into the scene, making them difficult for users to notice and for automated evaluation tools to detect.
@@
Detecting visual information manipulation attacks requires reasoning about both the virtual content and how it interacts with the real-world scene. Prior work on AR content quality assessment has primarily focused on evaluating visual properties of virtual elements, such as shadow consistency[3], lighting effect[24], and spatial alignment[28, 36, 55, 6] to determine how well the virtual content blends into the environment, or whether it affects users’ perception[53, 39, 18]. While important for visual coherence, these approaches fall short when it comes to semantic-level threats, where content may appear realistic and reasonable, yet alter the meaning of the scene. In such cases, virtual elements may adopt a consistent visual style while subtly introducing misinformation that leads to misunderstanding. Other work has focused more directly on safety concerns in AR, such as in the context of obstruction attacks[13, 32, 20, 59], where virtual content blocks real-world information and potentially leads to accidents or task failures. However, while obstruction attacks involve physical concealment, they do not capture scenarios where virtual content is non-obstructive or only partially overlaps with key information, yet still alters the semantic meaning of the scene. For example, in Fig.˜2, additional virtual guidance is inserted into the scene, directing the user toward an incorrect action without blocking any critical real-world element.
@@
Beyond AR-specific studies, there have been numerous studies in computer vision and multimodal machine learning that focused on scene understanding, including semantic segmentation[41, 15], image captioning[52, 27], and visual question answering (VQA)[5, 54]. These methods demonstrate impressive capabilities in extracting semantic meaning from complex visual inputs. However, they are typically developed and evaluated outside the AR context. As a result, these models’ ability to analyze the presence or potential impact of augmented elements remains underexplored. In AR scenarios, this ability to analyze virtual content is important: while users can only see the augmented view, a robust detection system should evaluate both the raw scene and its augmented counterpart to determine whether virtual content has introduced detrimental changes.
@@
These challenges call for new approaches to evaluate AR content, which go beyond visual realism or obstruction, and instead focus on the semantic impact of virtual elements. In particular, detecting whether virtual content manipulates real-world information requires understanding not just what objects are present, but how their meaning changes when combined. Recent advances in vision-language models (VLMs) with generative capabilities[43, 21, 4] have demonstrated strength in multimodal reasoning, including fine-grained scene interpretation and textual analysis. Moreover, VLMs have been proven effective in evaluating artificial content[23] and sensitive to virtual content in AR[19]. These capabilities make VLMs well-suited for assessing whether AR content introduces semantic inconsistencies or misleading information.
@@
We introduce AR-VIM, a dataset of 452 raw-AR video pairs collected in 202 different real-world scenes. The videos are annotated to reflect various attack types and semantic distortions. To validate the dataset labels, we obtained human annotations through a user-based labeling process, approved by Duke University Campus Institutional Review Board. The results demonstrate that our labeling aligns closely with human perception. The dataset is available on GitHub.111https://github.com/YM-Xiu/AR-VIM
We develop a multimodal semantic reasoning system, VIM-Sense, that combines VLMs with OCR-based text extraction to detect and analyze semantic-level information manipulation attacks in AR scenes. On AR-VIM, our system reaches an attack detection accuracy of 88.94% and a detection latency of 7.07 seconds. We also evaluate our system on a real-world Android AR application, achieving a detection latency of 7.17 seconds.
The remaining sections of this paper are organized as follows: Section˜1 reviews prior work, followed by Section˜2, which introduces our taxonomy of visual information manipulation attacks. Section˜3 describes the construction and annotation of the AR-VIM dataset. Section˜4 presents the design and implementation of our system, VIM-Sense. In Section˜5, we evaluate the VIM-Sense both on the AR-VIM dataset and in a real-world AR application. Finally, we discuss the limitations and directions for future work in Section˜6 and conclude the paper in Section˜7.
@@ 1.1 AR Virtual Content Assessment
Augmented reality research has long examined how virtual content can be designed and evaluated to ensure a pleasant user experience. A significant body of work focuses on assessing the visual properties of virtual elements to determine whether they blend seamlessly into the real world. Kim et al.[28] investigated artifacts of virtual content that arise when physics simulations are used to model interactions between real and virtual objects in AR. Li et al. [36] proposed a method for evaluating misalignment between real and virtual objects in a head-mounted AR assembly assistance system using depth map comparison. Golyadkin et al.[22] proposed a data synthesis and benchmarking framework for manga colorization, which incorporates image inpainting techniques to support consistent colorization of sequential pages and assess the inpainting quality. On the user perception side, Wang et al.[53] conducted a user study to investigate how texture density, luminance contrast, and color contrast of virtual objects affect depth perception in optical see-through AR. Liu et al.[39] considered visual confusion as a type of distortion affecting perceptual quality when both real and virtual contents contain complex textures. Duan et al.[18] proposed a framework for evaluating perceptual quality in AR scenes based on visual confusion theory. While these works provide insights into visual realism, perceptual quality, and alignment, they primarily focus on how well virtual content fits into the physical scene or how it affects users’ perception and do not analyze whether the content introduces semantic distortions or changes the meaning of the scene.
Aside from perceptual quality, some prior works have explored how virtual content can interfere with users’ access to real-world information. Cheng et al.[13] conduct a user study to investigate how people perceive and react to visual obstruction attacks in AR. Lebeck et al. proposed Arya[32, 33], which detects obstructions by using system sensors to identify predefined safety-critical real-world objects. It ensures that critical items are not obstructed by modifying virtual content based on predefined policies. In our prior work, we proposed VIDDAR[59, 58], which integrates generative AI models with object detection models to identify critical real-world objects and determine if a key object is obstructed or made to be visually misleading. These approaches take scene information and user safety into consideration when assessing the virtual content. However, they share a common assumption that safety threats in AR arise when important real-world objects are physically occluded. Consequently, they cannot handle more subtle cases of visual information manipulation, where virtual content does not fully block real-world objects but instead modifies or replaces their meaning.
@@ 1.2 Scene Understanding
Recently, numerous studies in computer vision have focused on scene understanding, which aims to extract semantic meaning from images through tasks such as semantic segmentation, image captioning, and visual question answering (VQA). Long et al.[41] proposed fully convolutional networks for semantic segmentation, which assigns a categorical label to each pixel in an image. Vinyals et al.[52] proposed an image captioning system that uses deep neural networks to generate natural language descriptions for the image. Antol et al.[5] proposed the task of VQA, which requires answering open-ended questions about images using natural language. These early works laid the foundation for a wide range of scene understanding tasks. More recently, the field has seen a shift toward foundation models. Kirillov et al.[29] proposed the Segment Anything Model (SAM), a foundation model for semantic image segmentation that supports prompt-based interaction. Generative models with visual modality, also known as VLMs, such as GPT-4V[43], Gemini[21], and Claude[4], have shown superior performance in some scene understanding benchmarks[12, 5, 2, 26]. Overall, recent methods have demonstrated strong capabilities in understanding scene content and identifying potentially harmful or misleading information. However, most of these studies were conducted and evaluated using standard, general-purpose vision-language benchmarks. As a result, their potential for understanding virtual content in AR, where it is essential to analyze both the real-world scene and its augmented counterpart, remains largely underexplored.
@@ 1.3 VLMs for Semantic Reasoning in AR Settings
As mentioned above, the ability of scene understanding methods to operate effectively in AR settings remains unclear. Recently, efforts have begun to explore the use of VLMs for reasoning about artificial content such as user/AI-generated imagery, as well as virtual content presented in AR, especially in the context of safety and usability. Guo et al.[23] proposed UGCG-GUARD, a content moderation system for detecting image-based promotions of unsafe user-generated content, leveraging VLMs to perform contextual detection. Huang et al.[25] discussed the visual hallucination of VLMs on AI-generated images. Srinidhi et al. proposed XaiR[49], a platform that integrates VLMs with XR headsets, which enables real-time semantic reasoning and AR content placement. Sharma et al. proposed OCTO+[46], a framework for automatic open-vocabulary object placement in AR, which leverages VLMs to reason about spatial semantics and place virtual objects in valid regions. Duan et al.[19] evaluated several state-of-the-art VLMs for understanding and describing AR-generated scenes, highlighting that VLMs excel at detecting and assessing obvious virtual content. Nonetheless, these work did not dive into the visual safety concerns in AR settings. In this work, we aim to fill this gap by combining VLM-based semantic reasoning with OCR-based textual grounding to detect visual information manipulation attacks in AR scenes.
@@ 2.1 Attack Format
Character Manipulation: Character manipulation refers to attacks where individual characters in a real-world scene are modified through virtual overlays. These changes may appear subtle (e.g., "3" to "8" or "O" to "Q"), but can significantly alter the meaning of textual information, such as room numbers, prices, or signs.
Denote as the ordered sequence of characters in , and as the ordered sequence of characters in . These two sequences have exactly the same length, which means the characters are only manipulated, without being erased or added. Therefore, the character in is manipulated if:
@@
Pattern Manipulation: Pattern manipulation refers to non-textual visual elements in the scene, such as colors, symbols, or directional cues, that are altered through virtual content in a way that changes the scene’s semantic meaning. Unlike character or phrase manipulation, which operates on textual information, pattern manipulation affects visual structures that convey meaning independently of text.
@@ 2.2 Attack Purpose
This condition indicates that the original information is no longer present in the AR image and has been replaced by another token that conveys a different semantic meaning. The substitution may occur in either the textual domain () or the visual domain (), as long as the replacement alters the perceived meaning of the original scene content.
@@
Extra Wrong Information: Extra wrong information attack is a case where new virtual content introduces semantic elements into the AR image that were not present in the raw image and that convey a misleading or incorrect meaning. Unlike information replacement or obfuscation, this type of attack does not interfere with existing real-world information but instead adds false signals that may confuse the user or alter the interpretation of the scene. An additional wrong information attack is said to occur if:
@@ 2.3 Attack Taxonomy Summary
Based on the definitions above, we categorize visual information manipulation attacks in AR by combining three attack formats with three attack purposes. This gives a conceptual space of possible attack types. However, not all combinations are applicable in real-world AR scenarios. In particular, we only consider character manipulation in the context of information replacement, since manipulating characters at the level of obfuscation or false addition is rare or ill-defined. For example, it is difficult to obfuscate a character without affecting its surrounding word or phrase, and adding an isolated character into the scene without meaningful context rarely conveys independent semantic intent. The resulting taxonomy consists of seven valid attack types, as shown in Table˜1. Each type corresponds to a unique way that virtual content can semantically alter the scene, whether through subtle text edits or high-level visual modifications. Some examples of the attacks discussed in the taxonomy are shown in Fig.˜3.
@@ 3.1 VIM Attack Design Methodology
Semantic Alteration with Minimal Virtual Content Insertion: For each scene, our objective was to introduce the minimal amount of virtual content required to alter the scene’s meaning. For example, in character replacement, we selected individual characters whose modification would flip the semantic interpretation (e.g., "F" to "E" and "1" to "7") while leaving the surrounding context unchanged. This ensures the manipulation is subtle yet capable of misleading the user.
Visual Style Coherence: To avoid triggering user attention, we ensured that the visual style of the virtual content closely matched the real-world context, which includes color style, lighting effect, and spatial alignment. This is especially important for information replacement and information obfuscation attacks, where the effectiveness of the manipulation relies on the user’s inability to recognize the virtual content. For extra wrong information attacks, we focused on adding semantically misleading but visually plausible content.
Attack-Non-Attack Pairing Strategy: To ensure controlled comparisons and reliable evaluation of VIM detection, we designed the corresponding non-attacked case for most of the attacked cases. For information obfuscation, we created a non-attacked version by repositioning the virtual content to a different part of the scene that does not obfuscate the key semantic elements in the real world. For information replacement and extra wrong information attacks, we generated visually similar virtual content in terms of color, style, and size, but with either misaligned placement or semantically innocuous meaning. For example, in Fig.˜3(b), the attack changes "traces of nuts" to "traces of milk", and the non-attacked version might replace "traces of nuts" with "nut ingredients" in the same font and location, preserving visual realism while avoiding deception.